OpenStudy (anonymous):

I have an algerbra question, I think I know the answer but looking at the question fully Im not sure...

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

yo :D

OpenStudy (anonymous):

The functions f(x) = –(x − 1)2 + 5 and g(x) = (x + 2)2 − 3 have been rewritten using the completing-the-square method. Is the vertex for each function a minimum or a maximum? Explain your reasoning for each function. I believe f(x) is a maximum and g(x) is a minimum

OpenStudy (anonymous):

from what I have read in my lesson that is what it leads me to believe, im not fully sure how to figure it out.

OpenStudy (anonymous):

@atreyu6s3x6 ? @Nnesha ?

Nnesha (nnesha):

okay so what's in ur notes what makes it max or min ?

OpenStudy (anonymous):

sorry i cant help D:

OpenStudy (anonymous):

ok thanks @atreyu6s3x6 @Nnesha in my lesson(I dont take notes... my handwriting is horrible) but it says "f(x) = a(x − h)2 + k" and if a is - it is maximum but if it is positive it is minimum.

Nnesha (nnesha):

right!! so in vertex form \[\rm y=-a(x-h)^2+k\] if a is negative u r going to get max point and if it's \[\rm y=a(x-h)^2+k\] u will get min point

Nnesha (nnesha):

so your answer is correct gO_Od job!

OpenStudy (anonymous):

:D thanks man, can I make a new post with another question so I can give you another medal?